# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v1

@@ Abstract
Reservoir computing is a machine learning framework characterized by its high computational ability and quick learning, making it well-suited for a wide range of applications for real-time computing. In this paper, we demonstrate that the memory capacity of a reservoir recurrent neural network scales sublinearly with the number of readout neurons. To elucidate this phenomenon, we develop a theoretical framework for analytically deriving memory capacity, attributing the decaying growth of memory capacity to neuronal correlations. In addition, numerical simulations reveal that once memory capacity becomes sublinear, increasing the number of readout neurons successively enables nonlinear processing at progressively higher polynomial orders. Furthermore, our theoretical framework suggests that neuronal correlations govern not only memory capacity but also the sequential growth of nonlinear computational capabilities. Our findings establish a foundation for designing scalable and cost-effective reservoir computing, providing novel insights into the interplay among neuronal correlations, linear memory, and nonlinear processing.
@@
Reservoir computing (RC) is a machine learning framework for efficiently training large scale recurrent neural networks (RNNs), referred to as ”reservoirs” in the context of RC [1, 2]. In contrast to conventional training methods such as backpropagation through time, RC optimizes only readout weights and leaves the remaining weights fixed, which enables quick and low-cost learning. In spite of its training simplicity, RC has high computational performance and a broad range of applications for processing time-series data [3]. RC is not restricted to RNNs, and a wide variety of dynamical systems can be utilized as reservoirs under appropriate conditions. Specifically, RC using a real physical system such as soft matter and optical systems, so called ”physical reservoir computing”, has been intensively studied in recent years [4].
@@
In a typical RC setting, the readout connections linking reservoir units to output units are sparse [4]. This means that the size of the reservoir ($N$) is substantially larger than the number of readout units ($L$) that connect to the output units. Empirically, the performance of RC improves as an increase in $L$. However, implementing a large number of readout connections can be resource-intensive, especially in physical RC implementations. Consequently, understanding the relationship between $L$ and the computational ability of a reservoir is essential for constructing cost-effective RC. In particular, performance in RC depends on both the reservoir’s memory capacity [5] and nonlinear processing capability [6], and the number of readout units must be carefully chosen to balance these two aspects according to the demands of a given task. Despite its practical importance, a systematic framework for determining an appropriate $L$ remains largely unexplored. In this study, we aim to address this gap by elucidating how memory capacity and nonlinear computational ability scale with $L$, with the particular focus on the role of neuronal correlations.
@@
Here, we demonstrate for the first time that the memory capacity of an RNN increases monotonically with $L$, though its growth rate progressively declines (Fig.1). This sublinear scaling of memory capacity cannot be explained by the previous theoretical works assuming a scaling regime where $L$ is of order one relative to $N$, i.e. $L\sim O(1)$ because the memory capacity in this regime is proportional to $L$  [7, 8]. To bridge the gap between the observation and theory, we develop a novel theory for analytically deriving memory capacity for $L\sim O(\sqrt{N})$. Using our theory, we demonstrate that the correlations between reservoir neurons, despite their very weak magnitude scaling as $O(1/\sqrt{N})$, play an important role in the declining growth rate of memory capacity. Furthermore, numerical simulations show that the ability to perform more complex nonlinear tasks emerges supralinearly and sequentially as $L$ increases. These findings indicate that the number of readout connections should be tailored to the specific memory and nonlinearity requirements of the task at hand.
@@
We investigate computational capacity of RC using a large random RNN as a reservoir (Fig.1(a)), which is known as Echo state network (ESN) [1], one of the canonical model for RC  [9, 10]. The state of $i$-th neuron at discrete time $t$ is described by the variable, $x_{i}(t)$. The activation function $\phi$ is assumed to be an odd saturated sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. The time evolution of a reservoir RNN is determined by the difference equation,
@@
where $N\gg 1$ indicates the total number of neurons. The recurrent weights, $J_{ij}$, and the input weights, $u_{i}$, are sampled i.i.d from Gaussian distributions, i.e. $J_{ij}\sim\mathcal{N}(0,g^{2}/N)$ and $u_{i}\sim\mathcal{N}(0,1)$. The parameter $g$ controls the strength of the recurrent weights. The input signal at time $t$, represented by $s(t)$, is Gaussian white noise with zero mean and variance $\sigma_{s}^{2}$. To take into account inherent neuronal noise, each neuron is subjected to independent Gaussian white noise, $\xi_{i}(t)$, with zero mean and variance $\sigma_{n}^{2}$.
The macroscopic dynamical behavior of the large random RNN described by Eq.(1) has been extensively studied by means of dynamical mean-field theory (DMFT) [11]. In the absence of the inputs and noise, i.e. $\sigma_{s}^{2}=\sigma_{n}^{2}=0$, the RNN exhibits phase transition from zero fixed-point to chaotic dynamics at $g=1$ in the large network size limit [11, 12]. In the presence of inputs ($\sigma_{s}^{2}>0$ or $\sigma_{n}^{2}>0$), chaotic variability is suppressed [7, 12, 13, 14, 8, 15].
The output of the reservoir RNN is defined as $\hat{z}(t)=\sum_{i=1}^{L}w_{i}x_{i}(t)$, where $L$ is the number of readout neurons. We assume that the readout neurons are sparse, i.e. $L\ll N$. According to the RC framework, the output weights, $\{w_{i}\}_{i=1}^{L}$, are optimized to minimize the time-averaged squared error between the output signals, $\{\hat{z}(t)\}_{t=1}^{T}$, and the desired signals, $\{z(t)\}_{t=1}^{T}$.
@@
$M_{d}\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(\hat{z}(t)-s(t-d))^{2}\rangle}{\langle s(t)^{2}\rangle},$ (2)
where the angular bracket denotes time averaging. A high value of $M_{d}$ indicates that the reservoir can accurately output its input signal from $d$ steps prior. It is shown that $0\leq M_{d}\leq 1$ holds true [5, 6]. Memory capacity is defined as the infinite sum of $M_{d}$, but it has been proven to be finite and bounded by the number of readout neurons, i.e. $0\leq MC\leq L$  [5, 6]. In particular, for RC with a linear activation function operating in noise-free conditions ($\sigma_{n}^{2}=0$), memory capacity is equivalent to $L$ independently of $g$ and $\sigma_{s}^{2}$  [5].
@@
By simple calculation, $M_{d}$ defined in Eq.(2) can be expressed as
@@
We noticed that from the scale assumption Eq.(3), the diagonal entries of $C$ are $O(1)$, whereas the non-diagonal ones are $O(1/\sqrt{N})$ (See Appendix.A). Hence, the non-diagonal elements are much smaller than the diagonal ones, which motivates us to perform Neumann series expansion,
@@
In the large network size limit, we can assume self-averaging for $M_{d}$; $\lim_{N\to\infty}M_{d}=\lim_{N\to\infty}[M_{d}]$, where the square bracket denotes the average over network realizations, known as quenched average [17]. Consequently, substituting Eq.(5) into Eq.(4) and taking quenched average of $M_{d}$, we analytically obtain memory capacity:
@@
We have confirmed that the analytical values of $MC$ are consistent with the numerical simulations, exhibiting a sublinear function of $L$, for $L$ below a threshold (Fig.2). When $L$ exceeds the threshold, the analytical values rapidly deviate from the simulated ones and eventually diverge (dashed lines in Fig.2). This is because Neumann series expansion, Eq.(5), does not converge as the spectrum norm $\|\tilde{C}D^{-1}\|$ exceeds one for large $L$. To evaluate the upper bound of $L$ ensuring convergence of the series expansion, it is enough to evaluate $\|\tilde{C}D^{-1}\|$, but this is a challenging task. Instead, we derive the sufficient condition for the expansion series to converge (see Appendix.B.1), and the upper values of $L$ for this sufficient condition are indicated by star marks in Fig.2.
@@
Having obtained the analytical solution of memory capacity, we move on to elucidate the reason why memory capacity grows sublinearly with respect to $L$. To quantify the extent to which the growth rate of memory capacity decays, we define a decay rate,
@@
It is noteworthy that the $n$-th term corresponds exactly to the $n$-th term in the Neumann series expansion, Eq.(5). Therefore, under the assumption of vanishing neuronal correlations ($\tilde{C}=O$), the summation terms in Eq.(Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks) vanish, resulting in $r(L)=1$. This implies that higher neuronal correlations contribute to the faster decay of $r(L)$ through the contributions of the summation terms in Eq.(Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks). In fact, we have confirmed that $r(L)$ decays more rapidly for hyperparameters that lead to higher neuronal correlation, such as larger input signals ($\tilde{\sigma}_{s}^{2}$), smaller recurrent weights ($g$), and lower noise intensity ($\sigma_{n}^{2}$), as depicted in Fig.6. Here, it is interesting to note that, while the slope of $MC(L)$ at the origin, which is proportional to $MC(1)$, reaches its maximum when the reservoir RNN operates near the edge of chaos [7, 8], the extent of the decay of $r(L)$ varies monotonically with the hyperparameters, irrespective of whether the RNN is near the edge of chaos.
@@
Information processing capacity (IPC) introduced by Dambre et al. [6] offers a task-independent metric for evaluating the performance of RC, enabling a comprehensive assessment of both linear and non-linear computational capabilities. Since IPC theory is somewhat complicated, we provide detailed explanations in Appendix.C.1. Put simply, the IPC for degree $D$, denoted by $IPC_{D}$, represents the reservoir’s ability to approximate $D$-th order polynomial functions of Gaussian white noise input signals. The IPC has two key properties. First, by definition, memory capacity is exactly equivalent to $IPC_{1}$, and thus, $IPC_{D}$ for $D\geq 2$ represents non-linear computational ability. Second, it has been proven [6] that the total IPC, $IPC_{\rm total}\equiv\sum_{D=1}^{\infty}IPC_{D}$, equals to the number of readout units, $L$, provided that the reservoir satisfies echo state property (ESP). Here, the ESP implies that the reservoir’s state is uniquely determined solely by input signals regardless of reservoir’s initial state [1].
Fig.4 illustrates how the values of $IPC_{D}$ for our RNN model (Eq.(1)) change with increasing $L$ across various hyperparameter combinations ($g$, $\sigma_{s}$, and $\sigma_{n}$), as obtained through numerical simulations. Note that, due to the symmetry of the model, $IPC_{D}$ values for even-degree $D$ are identically zero. As shown, a decline in the growth of the $IPC_{1}$ (equivalent to memory capacity) is accompanied by an increase in the non-linear computational ability represented by $IPC_{D}$ for $D\geq 2$, reflecting memory-nonlinearity trade-off.
Our novel finding is that, as $L$ increases, successively higher-degree IPCs are sequentially activated from lower to higher orders, accompanied by supralinear rises in the reservoir’s nonlinear computational capabilities. This phenomenon is observed not only in an RNN satisfying the echo state property (Fig.4(a)), but also in a chaotic RNN (Fig.4(b)) and in an RNN subject to neuronal noise (Fig.4(c)), where the echo state property breaks down. Additionally, the same finding is observed for an RNN employing the ReLU activation function, although our theoretical framework for deriving memory capacity is not applicable to the ReLU function (see Appendix.C.2). As analytical derivation of higher-degree IPC is beyond our capability, the detailed mechanism underlying these observations remains unclear. However, it is plausible that the neuronal correlation plays a pivotal role, analogous to its influence on memory capacity. This hypothesis is supported by speculation that each $IPC_{D}$ would exhibit linear scaling with $L$ if the effects of neuronal correlations were neglected, which contrasts with the observed supralinear and sequential emergence (see Appendix.C.3).
@@
In the present study, we have investigated the relationship between the number of readout neurons and the computational capacity for large random RNN reservoirs. Through analytical and numerical approaches, we have demonstrated that the memory capacity grows sublinearly with $L$, and that nonlinear computational capabilities emerge incrementally with $L$, enabling the reservoir to handle increasingly complex tasks. These findings suggest that the number of readout connections should be optimized according to the specific memory and nonlinearity demands of the task, which can be evaluated using the framework proposed by [21], thereby facilitating the design of cost-effective RC.
Our findings extend beyond the scope of RC. First, in neuroscience, while the influence of neuronal correlations on information processing has been extensively studied in the context of neural manifolds [22, 23, 24], the mammalian cortex in the awake state is known to operate in an asynchronous regime characterized by very weak neuronal correlations [25, 26, 27], similar to those observed in our large random RNNs. Our study offers a novel perspective on how weak neuronal correlations influence neural computation. Second, in deep learning theory, it has been established that neural networks trained via gradient-based methods can learn tasks in the lazy regime, where their weights undergo only infinitesimal changes from their initial random values [28, 29, 30]. Our insights may provide valuable understanding of the computational capacity of RNNs trained in the lazy regime.

### v2

@@ Abstract
Reservoir computing is a powerful framework for real-time information processing, characterized by its high computational ability and quick learning, with applications ranging from machine learning to biological systems. In this paper, we demonstrate that the memory capacity of a reservoir recurrent neural network scales sublinearly with the number of readout neurons. To elucidate this phenomenon, we develop a theoretical framework for analytically deriving memory capacity, attributing the decaying growth of memory capacity to neuronal correlations. In addition, numerical simulations reveal that once memory capacity becomes sublinear, increasing the number of readout neurons successively enables nonlinear processing at progressively higher polynomial orders. Furthermore, our theoretical framework suggests that neuronal correlations govern not only memory capacity but also the sequential growth of nonlinear computational capabilities. Our findings establish a foundation for designing scalable and cost-effective reservoir computing, providing novel insights into the interplay among neuronal correlations, linear memory, and nonlinear processing.
I Introduction
Over the past decade, deep learning has advanced significantly, with generative AI in particular beginning to transform many aspects of daily life. Although deep neural networks possess the capacity to represent highly complex functions, they typically require substantial computational resources, large volumes of training data, and long training times. These high computational demands pose a major obstacle to applying deep learning in real-time computing on edge-computing devices [1].
@@ I Introduction
Reservoir computing (RC) is a machine learning framework for efficiently training large scale recurrent neural networks (RNNs), referred to as ”reservoirs” in the context of RC [2, 3]. In contrast to conventional training methods such as backpropagation through time (BPTT), RC optimizes only readout weights and leaves the remaining weights fixed, which enables quick and low-cost learning. In spite of its training simplicity, RC has high computational performance and a broad range of applications for processing time-series data [4]. RC is not restricted to RNNs, and a wide variety of dynamical systems can be utilized as reservoirs under appropriate conditions. In particular, RC using a real physical system such as soft matter and optical systems, so called ”physical reservoir computing”, has been intensively studied in recent years [5].
@@
In a typical RC setting, the readout connections linking reservoir units to output units are sparse [5]. This means that the size of the reservoir ($N$) is substantially larger than the number of readout units ($L$) that connect to the output units. Empirically, the performance of RC improves as an increase in $L$. However, implementing a large number of readout connections can be resource-intensive, especially in physical RC implementations. Consequently, understanding the relationship between $L$ and the computational ability of a reservoir is essential for constructing cost-effective RC. In particular, performance in RC depends on both the reservoir’s memory capacity [6] and nonlinear processing capability [7], and the number of readout units must be carefully chosen to balance these two aspects according to the demands of a given task. Despite its practical importance, a practical principle for determining an appropriate $L$ remains largely unexplored. In this study, we aim to address this gap by elucidating how memory capacity and nonlinear computational ability scale with $L$, with the particular focus on the role of neuronal correlations.
@@
We demonstrate, for the first time, that the memory capacity of an RNN increases monotonically with $L$, although its growth rate progressively declines (Fig.1). This sublinear scaling of memory capacity cannot be explained by the previous theoretical works assuming a scaling regime where $L$ is of order one relative to $N$, i.e. $L\sim O(1)$ [8, 9], because the memory capacity in this regime is proportional to $L$, as will be shown later. To bridge the gap between the observation and theory, we develop a novel theory for analytically deriving memory capacity for $L\sim O(\sqrt{N})$. Using our theoretical framework, we demonstrate that the correlations between reservoir neurons, despite their very weak magnitude scaling as $O(1/\sqrt{N})$, play an important role in the declining growth rate of memory capacity. Furthermore, numerical simulations show that the ability to perform more complex nonlinear tasks emerges supralinearly and sequentially as $L$ increases. These findings indicate that the number of readout connections should be tailored to the specific memory and nonlinearity requirements of the task at hand.
II Model setup
@@
We investigate computational capacity of RC using a large random RNN as a reservoir (Fig.1(a)), which is known as Echo state network (ESN) [2], one of the canonical model for RC  [10, 11]. The state of $i$-th neuron at discrete time $t$ is described by the variable, $x_{i}(t)$. The activation function $\phi$ is assumed to be an odd saturating sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. The time evolution of a reservoir RNN is determined by the difference equation,
@@
where $N\gg 1$ indicates the total number of neurons. The recurrent weights, $J_{ij}$, and the input weights, $u_{i}$, are sampled i.i.d from Gaussian distributions, i.e. $J_{ij}\sim\mathcal{N}(0,g^{2}/N)$ and $u_{i}\sim\mathcal{N}(0,1)$. The parameter $g$ controls the strength of the recurrent weights. The input signal at time $t$, represented by $s(t)$, is Gaussian white noise with zero mean and variance $\sigma_{s}^{2}$. To take into account inherent neuronal noise, each neuron is subjected to independent Gaussian white noise, $\xi_{i}(t)$, with zero mean and variance $\sigma_{n}^{2}$. Notably, the input signal – shared across all neurons – induces correlations among reservoir neurons, whereas the independent neuronal noise acts to diminish them.
The macroscopic dynamical behavior of the large random RNN described by Eq.(1) has been extensively studied by means of dynamical mean-field theory (DMFT) [12]. In the absence of the inputs and noise, i.e. $\sigma_{s}^{2}=\sigma_{n}^{2}=0$, the RNN exhibits phase transition from zero fixed-point to chaotic dynamics at $g=1$ in the large network size limit [12, 13]. In the presence of inputs ($\sigma_{s}^{2}>0$ or $\sigma_{n}^{2}>0$), chaotic variability is typically suppressed [8, 13, 14, 15, 9, 16], although it has recently been shown that in RNNs incorporating gating mechanisms similar to those in gated recurrent units (GRUs), external inputs can instead induce chaos [17].
The output of the reservoir RNN is defined as $z(t)=\sum_{i=1}^{L}w_{i}x_{i}(t)$, where $L$ is the number of readout neurons. We assume that the readout neurons are sparse, i.e. $L\ll N$. According to the RC framework, the output weights, $\{w_{i}\}_{i=1}^{L}$, are optimized to minimize the time-averaged squared error between the output signals, $\{z(t)\}_{t=1}^{T}$, and the desired signals, $\{f(t)\}_{t=1}^{T}$.
III Memory capacity
@@
$M_{d}\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-s(t-d))^{2}\rangle}{\langle s(t)^{2}\rangle},$ (2)
and the angular bracket denotes time averaging. A high value of $M_{d}$ indicates that the reservoir can accurately output its input signal from $d$ steps prior. It is shown that $0\leq M_{d}\leq 1$ holds true [6, 7]. Memory capacity is defined as the infinite sum of $M_{d}$, but it has been proven to be finite and bounded by the number of readout neurons, i.e. $0\leq MC\leq L$  [6, 7]. In particular, for RC with a linear activation function operating in noise-free conditions ($\sigma_{n}^{2}=0$), memory capacity is equivalent to $L$ independently of $g$ and $\sigma_{s}^{2}$  [6].
@@ IV Theoretical analysis of memory capacity
By simple calculation, $M_{d}$, defined in Eq.(2), can be expressed as
@@
We noticed that from the scale assumption Eq.(3), the diagonal entries of $C$ are $O(1)$, whereas the non-diagonal ones are $O(1/\sqrt{N})$ (See Appendix.A.2). Hence, the non-diagonal elements are much smaller than the diagonal ones, which motivates us to perform Neumann series expansion,
@@
In the large network size limit, we can assume self-averaging for $M_{d}$, $\lim_{N\to\infty}M_{d}=\lim_{N\to\infty}[M_{d}]$, where the square bracket denotes the average over network realizations, known as quenched average [19]. Consequently, substituting Eq.(5) into Eq.(4) and taking quenched average of $M_{d}$, we analytically obtain memory capacity:
@@
In deriving this result, it is necessary to calculate quenched averages of quantities involving neuronal correlations, such as
$\left[\langle s(t-d)x_{i}(t)\rangle\langle x_{i}(t)x_{j}(t)\rangle\langle s(t-d)x_{j}(t)\rangle\right]\ \ (i\neq j).$ (8)
However, these quantities cannot be evaluated using DMFT because it reduces the original $N$-body system to an effective single-body description, thereby neglecting inter-neuronal correlations. To overcome this limitation, we adopt the dynamical cavity method, a technique originally introduced by Clark et.al [20] from statistical physics to the analysis of neural networks. A detailed description of this theory is provided in Appendix.A.1 and A.2, followed by the full derivation of Eq.(6) in Appendix.B.1. In the following, we use the shorthand notation, $\langle f(x)\rangle_{*}\equiv\langle f(x)\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}$.
@@
$MC=\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(\frac{g^{2}}{1+\frac{\pi}{2}[\langle x_{i}^{2}\rangle]}\right)^{d}\right\}^{n+1},$
where $[\langle x_{i}^{2}\rangle]$ is determined by solving
$[\langle x_{i}^{2}\rangle]=\sigma_{n}^{2}+g^{2}\left(-1+\frac{4}{\pi}\arctan\sqrt{1+\pi[\langle x_{i}^{2}\rangle]}\right).$
We have confirmed that the analytical values of $MC$ are consistent with the numerical simulations, exhibiting a sublinear function of $L$, for $L$ below a threshold (Fig.2). When $L$ exceeds the threshold, the analytical values rapidly deviate from the simulated ones and eventually diverge (dashed lines in Fig.2). This is because Neumann series expansion, Eq.(5), does not converge as the spectrum norm $\|\tilde{C}D^{-1}\|$ exceeds one for large $L$. To evaluate the upper bound of $L$ ensuring convergence of the series expansion, it is enough to evaluate $\|\tilde{C}D^{-1}\|$, but this is a challenging task. Instead, we derive the sufficient condition for the expansion series to converge (see Appendix.B.1), and the corresponding upper bounds of $L$ satisfying this condition are indicated by star marks in Fig.2.
@@
V Neuronal correlations shape the sublinear scaling of memory capacity
Having obtained the analytical solution of memory capacity, we now investigate the underlying reason why memory capacity grows sublinearly with respect to $L$. To quantify the extent to which the growth rate of memory capacity decays, we define a decay rate,
@@ V Neuronal correlations shape the sublinear scaling of memory capacity
It is noteworthy that the $n$-th term corresponds exactly to the $n$-th term in the Neumann series expansion, Eq.(5). Therefore, under the assumption of vanishing neuronal correlations ($\tilde{C}=O$), the summation terms in Eq.(V) vanish, resulting in $r(L)=1$. This implies that higher neuronal correlations contribute to the faster decay of $r(L)$ through the contributions of the summation terms in Eq.(V). In fact, we have confirmed that $r(L)$ decays more rapidly for hyperparameters that lead to higher neuronal correlation, such as larger input signals ($\tilde{\sigma}_{s}^{2}$), smaller recurrent weights ($g$), and lower noise intensity ($\sigma_{n}^{2}$), as depicted in Fig.3. Notably, while the slope of the memory capacity curve at the origin, which is proportional to $MC(1)$, reaches its maximum when the reservoir RNN operates near the edge of chaos [8, 9], the extent of the decay of $r(L)$ varies monotonically with the hyperparameters $g$, $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$, irrespective of whether the RNN is near the edge of chaos. These results suggest that instead of proximity to the edge of chaos, neuronal correlation is a novel mechanism determining the computational ability of RC.
[…]
@@ VI Scaling behavior of nonlinear computational capability
Information processing capacity (IPC) introduced by Dambre et al. [7] offers a task-independent metric for evaluating the performance of RC, enabling a comprehensive assessment of both linear and non-linear computational capabilities. Since IPC theory is somewhat complicated, we provide detailed explanations in Appendix.C.1. Put simply, the IPC for degree $D$, denoted by $IPC_{D}$, represents the reservoir’s ability to approximate $D$-th order polynomial functions of past white noise input signals. The IPC has two key properties. First, by definition, memory capacity is exactly equivalent to $IPC_{1}$, and thus, $IPC_{D}$ for $D\geq 2$ represents non-linear computational ability. Second, it has been proven [7] that the total IPC, defined as $IPC_{\rm total}\equiv\sum_{D=1}^{\infty}IPC_{D}$, equals to the number of readout units, $L$, provided that the reservoir satisfies the echo state property, a necessary condition for a reservoir to be usable for RC. Here, the echo state property implies that the reservoir’s state is uniquely determined solely by input signals regardless of reservoir’s initial state [2].
Fig.5 illustrates how the values of $IPC_{D}$ for our RNN model (Eq.(1)) change with increasing $L$ across various hyperparameter combinations ($g$, $\sigma_{s}^{2}$, and $\sigma_{n}^{2}$), as obtained through numerical simulations. Note that, due to the symmetry of the model, $IPC_{D}$ values for even-degree $D$ are identically zero. As shown, a decline in the growth of the $IPC_{1}$ (equivalent to memory capacity) is accompanied by an increase in the non-linear computational ability represented by $IPC_{D}$ for $D\geq 2$, reflecting memory-nonlinearity trade-off.
Our novel finding is that, as $L$ increases, successively higher-degree IPCs are sequentially activated from lower to higher orders, accompanied by supralinear rises in the reservoir’s nonlinear computational capabilities. This phenomenon is observed not only in an RNN satisfying the echo state property (Fig.5(a)), but also in a chaotic RNN (Fig.5(b)) and in an RNN subject to neuronal noise (Fig.5(c)), where the echo state property breaks down. Additionally, the same finding is observed for an RNN employing the ReLU activation function (Fig.8), where our theoretical framework for deriving memory capacity is not applicable (see Appendix.C.2).
As analytical derivation of higher-degree IPC is beyond our capability, the detailed mechanism underlying these observations remains unclear. However, it is plausible that the neuronal correlation plays a pivotal role, analogous to its influence on memory capacity. This hypothesis is supported by speculation that each $IPC_{D}$ would exhibit linear scaling with $L$ if the effects of neuronal correlations were neglected, which contrasts with the observed supralinear and sequential emergence of $IPC_{D}$ (see Appendix.C.3).
@@
VII Discussion
In the present study, we have investigated the relationship between the number of readout neurons and the computational capacity for large random RNN reservoirs. Through analytical and numerical approaches, we have demonstrated that the memory capacity grows sublinearly with $L$, and that nonlinear computational capabilities emerge incrementally with $L$, enabling the reservoir to handle increasingly complex tasks. The specific memory and nonlinearity requirements of a given task can be assessed using the framework proposed by [23]. In this context, our results provide a practical principle for designing cost-effective RC systems: while increasing $L$ raises computational and structural costs and yields diminishing returns for memory, it becomes essential when tasks demand higher-order nonlinear computations.
Our findings have implications that extend beyond the framework of RC. First, in neuroscience, the impacts of neuronal correlations on function have been extensively investigated from the perspective of neural manifold [24, 25, 26]. A neuronal network with a high level of correlations occupies a restricted region of $N$-dimensional phase space, whose linear dimension is quantified by effective dimension [27]. In the context of population coding, a well-established trade-off exists between representation richness and robustness: high-dimensional (low-correlation) activity enables complex and diverse information encoding, while low-dimensional (high-correlation) activity supports redundant encoding and noise robustness [25, 26]. However, the mammalian cortex in the awake state is known to operate in an asynchronous regime characterized by near-zero neuronal correlations [28, 29, 30], akin to those observed in our large random RNNs. In addition, these prior studies have only focused on the implications of neuronal correlations for static information processing, i.e., mapping static inputs to outputs. Our findings suggest a novel role for weak correlations in dynamic information processing, offering a new hypothesis about their functional significance in the brain.
Second, in the field of deep learning theory, it has been established that neural networks trained via gradient-based methods, such as BPTT, can learn tasks both in the lazy regime, where their weights undergo only infinitesimal changes from their initial random values [31, 32], and in the rich regime, where weights change substantially to extract task-relevant features [33, 34, 35]. Our insights may provide valuable understanding of the computational capacity of RNNs trained in the lazy regime, as their recurrent weights remain nearly identical to those of untrained, reservoir-style networks.